# Dynamics of holographic steady flows near a first-order phase transition

### v1

@@ Abstract
We investigate the physical properties of steady flows in a holographic first-order phase transition model, from the thermodynamics at equilibrium to the real-time dynamics far from equilibrium. Through spinodal decomposition and condensation nuclei, the phase-separated state with non-zero momentum is obtained, respectively. In such a geometric configuration, in addition to energy density, there is also a gap between coexisting phases for momentum density or longitudinal pressure, which is found to be characterized by flow velocity and latent heat. Furthermore, by introducing an inhomogeneous scalar external source to simulate a fixed obstacle, the dynamical response of momentum loss of the moving system is revealed. In particular, using the phase-separated state with uniform flow velocity as the background, we find that the moving high-energy phase exhibits four characteristic dynamical behaviors when interacting with an obstacle, rebounding, pinning, passing and splitting, depending on its velocity and the strength of the obstacle.
Keywords:
Black Holes, Gauge/Gravity Correspondence, Holography and Hydrodynamics, Holography and quark-gluon plasmas
1 Introduction
@@
The dynamics of strongly coupled quantum fluids is of great research significance in particle physics experiments and astronomical observations, such as the transport properties of the quark gluon-plasma (QGP) Schafer:2009dj and the merger process of binary neutron stars Alford:2017rxf . The challenge is to construct a theoretical framework to describe the real-time evolution of strongly interacting matter. Fortunately, by dualizing a quantum many-body system to a classical gravitational entity with an extra dimension, holography Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj ; Witten:1998zw converts such a problem into solving a system of time-dependent gravitational equations, which is always amenable to the numerical relativity Chesler:2013lia .
@@
A breakthrough related to holographic fluids is the description of dynamics near a thermal first-order phase transition. Initially, in order to mimic the equation of state of quantum chromodynamics (QCD), the holographic first-order phase transition model was constructed bottom-up Gubser:2008yx ; Gubser:2008ny , involving a real self-interacting scalar field with an external source. Subsequently, the dynamical stability of the thermal phases in equilibrium was revealed through linear perturbation theory Janik:2015iry ; Janik:2016btb . The results indicate the existence of a branch of hydrodynamic modes that dominates the linear stability of the system. For the spinodal region with negative specific heat, the hydrodynamic modes become dynamically unstable in the long wavelength region, similar to Gregory-Laflamme instability Gregory:1993vy ; Gregory:1994bj . Going a step further, the real-time dynamics of the spinodal instability was numerically simulated at the nonlinear level Attems:2017ezz ; Janik:2017ykj ; Bellantuono:2019wbn ; Attems:2019yqn ; Attems:2020qkg , revealing the second-order hydrodynamic description of the dynamical process and the phase-separated configuration of the final state. On the other hand, for the thermodynamically metastable region adjacent to the spinodal region, the occurrence of a dynamical transition requires the system to cross an unstable critical state Bea:2020ees ; Bea:2022mfb ; Chen:2022cwi , in which a critical phenomenon analogous to that in gravitational collapse Choptuik:1996yg ; Liebling:1996dx ; Bizon:1998kq ; Gundlach:2007gc occurs, indicating a nonlinear instability. Furthermore, through holographic quench, the nonlinear instability of the phase-separated state was also revealed Chen:2022tfy , achieving a bidirectional dynamical transition between homogeneous and inhomogeneous configurations. At this point, the theoretical framework on the dynamical stability of the holographic first-order phase transition system has been roughly completed.
@@
Beyond stability analysis, the dynamic properties of holographic fluids near phase transitions and their applications have also been extensively investigated. An interesting research topic is the simulation of cosmological phase transitions Kibble:1980mv ; Hindmarsh:2020hop , with the resulting gravitational wave spectrum potentially observable to detectors Ares:2020lbt ; Ares:2021nap . For the case of cosmological phase transitions proceed via bubble nucleation, the dynamic description of the domain wall connecting the nucleated bubble of a stable phase and the overcooled medium was successfully revealed Bea:2021zsu ; Janik:2021jbq ; Janik:2022wsx , in which the velocity of the domain wall, a crucial parameter on which the gravitational wave spectrum depends, is quantitatively characterized. On the other hand, if the Universe supercools along the metastable branch and enters the spinodal region, there will be a cosmological phase transition induced by the exponential growth of unstable modes, resulting in a qualitatively different gravitational wave spectrum Bea:2021zol . Another important application lies in the QGP produced in relativistic heavy ion collisions Casalderrey-Solana:2011dxg , which are viewed as collisions of gravitational shock waves in the dual description Chesler:2010bi ; Casalderrey-Solana:2013aba ; Casalderrey-Solana:2013sxa ; Chesler:2015wra ; Chesler:2015bba ; Chesler:2016ceu . If QCD possesses a critical point, then the far-from-equilibrium dynamics of QGP across a phase transition can be effectively captured by a holographic model with a phase transition Attems:2016tby ; Attems:2018gou ; Bea:2021ieq . Other related content includes holographic turbulence Adams:2012pj ; Adams:2013vsa , holographic superfluid Herzog:2008he ; Bhaseen:2012gg , etc.
@@
Inspired by the fact that fluids in reality generally have a flow velocity, we will study the real-time dynamics of the holographic steady flow, which is dual to the boosted black brane geometry on the gravity side Bhattacharyya:2007vjd ; Ecker:2021ukv , near a thermal first-order phase transition. The plan of the paper is as follows. In the following section 2, we will introduce a holographic first-order phase transition model through the thermodynamic properties of the equilibrium system. In section 3, we will study the dynamical transitions of holographic steady flows through fully nonlinear numerical simulations, revealing spontaneous and critical dynamical processes, which correspond to unstable and metastable initial states, respectively. Similarly, the final state is a two-phase coexisting geometric configuration. The difference is that the high-energy phase moves in the low-energy bath at a uniform velocity.
@@
A natural question is whether there is a damping mechanism that slows down the moving domain. We will provide such a mechanism in section 4 by introducing an inhomogeneous external source to the scalar field. From the Ward-Takahashi identity, such a local inhomogeneous scalar source can effectively induce changes in the momentum of the system, thereby damping the motion of a steady flow or a moving domain, which is verified by the dynamical results. In particular, we reveal four characteristic dynamical behaviors of the moving domain under interaction with the local potential barrier: rebounding, pinning, passing through, and splitting, depending on the flow velocity of the moving domain and the strength of the potential barrier. Finally, we end this paper with a conclusion and outlook in section 5.
2 Holographic model
@@
In this section, we will introduce a holographic model with a thermal first-order phase transition from equations of motion, to Ward-Takahashi identity, to thermodynamic phase diagrams involving static and steady states.
2.1 Gravity setup
To avoid logarithmic terms in the near-boundary expansion of the geometry, which would destroy the stability of numerical calculations, we consider a three-dimensional field theory with four-dimensional gravitational duality, with the following renormalized action Gibbons:1976ue ; Bianchi:2001kw ; Elvang:2016tzz
@@ 2.1 Gravity setup
According to the AdS/CFT correspondence, such a gravity system is dual to a boundary conformal field theory deformed with a scalar operator.
@@
where all metric fields are functions of $(t,x,r)$, and the translational invariance in the $y$-direction is preserved. In this case, the non-zero components of the energy-momentum tensor (6a) are $\left(\epsilon,\eta,p_{L},p_{T}\right)=\left(-T^{t}_{t},-T^{t}_{x},T^{x}_{x},T^{y}_{y}\right)$, which represent the energy density, momentum density, the longitudinal and transverse pressures of the system, respectively. The non-trivial components of the Ward-Takahashi identity (8) reduce to
@@
Based on the second relationship above (10b), we expect that the local inhomogeneous scalar source can provide a damping mechanism for the moving system to evolve from a steady state to a static state.
2.2 Phase diagram
@@
In order to obtain the steady state with non-zero momentum density, we need to abandon such an assumption and retain all components of the metric ansatz (9). The system at this time has two independent physical degrees of freedom. Note that in this section we are concerned with the time-independent homogeneous solution, where the field functions depend only on the coordinate $r$, and set the scalar source to $\phi_{0}=1$. After obtaining the geometric configuration and setting the effective Newton constant $\kappa^{2}_{4}=1$, one can extract the temperature and entropy density of the system as follows
@@
where $r_{h}$ represents the location of the event horizon. The free energy density satisfies the following thermodynamic relation
@@
Figure 1 shows the phase diagrams under different flow velocities, which are related through the usual boost transformation. For a steady state with flow velocity $v$, after applying the Lorentz transformation
@@
with a boundary one-form $\omega=\gamma\left(-dt^{\prime}+vdx^{\prime}\right)$. Such a metric form describes a comoving fluid system relative to an observer whose $3$-velocity components equal $\omega^{i}$ Chesler:2013lia . To this moving observer, the fluid is at rest. Note that in the frame (9), the boundary one-form is $-dt$, indicating a stationary observer. Our numerical results show that for any observer, the relatively stationary system has the same phase structure, that is, the case of $v=0$ in figure 1. In other words, in the frame (14), the phase structure of the system does not depend on the parameter $v$.
@@
Quantitatively, figure 2 shows the phase transition temperature and latent heat of the first-order phase transition as a function of the flow velocity. The monotonically decreasing curve in figure 2 indicates that a steady flow with a faster flow velocity has a lower phase transition temperature. Such a result shows that for a fast-flowing fluid, the first-order phase transition is difficult to occur. An interesting question is what happens when the flow velocity approaches the speed of light. Unfortunately, due to numerical problems, we cannot obtain the geometric configuration for the case $v>0.8$, leaving this question unresolved and awaiting further exploration in the further. As shown in figure 2, the latent heat of phase transition increases monotonically with the increase of flow velocity, indicating that the faster the fluid flows, the more energy needs to be absorbed or released to transform between different thermal phases. On the other hand, from relationship (17), the increasing latent heat ensures that the pressure difference between the coexisting phases increases with the increase of flow velocity, providing sufficient driving force for the movement of the domain wall.
3 Phase separation
@@
3.1 Spinodal decomposition
For the spinodal region where the specific heat is negative, the thermodynamic instability of the system induces a dynamical instability Janik:2015iry ; Janik:2016btb . At the linear level, there is a branch of hydrodynamic modes where the imaginary part of the frequency is positive in the long wavelength region. Under arbitrarily small perturbations, such unstable modes will grow exponentially and push the system away from the initial equilibrium state, inducing a dynamical transition.
In order to reveal such a dynamical process in the case of steady flow, we select a steady state in the spinodal region as the initial data and then apply a small perturbation. Figure 3 shows the real-time dynamics of the energy density of the system, from which it can be seen that a dynamical transition occurs that destroys the spatial translational symmetry. The entire dynamical process can be roughly divided into three stages. Initially, several energy peaks are spontaneously excited from the homogeneous geometric configuration and move along the flow direction of the background. Subsequently, these energy peaks merge together to form a moving domain, accompanied by fluctuations in energy density. Finally, as the energy fluctuations decay, the system gradually settles down to a phase-separated state with non-zero momentum. During this evolution, from the Ward-Takahashi identity (10), the total energy and total momentum of the system are conserved.
The final state of the evolution consists of coexisting phases at the phase transition temperature, which occupy different spatial regions and are connected by domain walls. The conservation of total momentum results in the system not being at rest, but flowing steadily at a uniform velocity as in the initial state. Since the momentum density of the low-energy phase is close to zero, as shown in figure 1, such a geometric configuration can be approximated as a dynamic structure of a moving domain of the high-energy phase in the background of a static low-energy bath. Note that the velocity of the moving domain in the final state is $v=0.37$, compared with the velocity $v=0.4$ of the initial steady flow, indicating that the phase transition reduces the flow velocity of the system. For such a time-dependent inhomogeneous steady state with uniform flow velocity $v$, there exists a Killing vector
@@ 3.1 Spinodal decomposition
3.2 Condensation nucleus
For the thermodynamically metastable region adjacent to the spinodal region, the system is dynamically stable at the linear level due to the absence of unstable modes, indicating that small perturbations will be dissipated and are not sufficient to change the state of the system. However, inspired by the fact that condensation nuclei induce phase transitions of supercooled water in reality, the thermodynamically metastable state is verified to suffer from a type of nonlinear instability Bea:2020ees ; Bea:2022mfb ; Chen:2022cwi . In this case, the system is in a local ground state. The occurrence of a phase transition requires a disturbance of sufficient intensity to push the system across a dynamical barrier and thereby evolve to another local ground state. Near the threshold of the disturbance intensity, a critical state will appear.

### v3

@@ Abstract
We investigate the physical properties of steady flows in a holographic first-order phase transition model, extending from the thermodynamics at equilibrium to the real-time dynamics far from equilibrium. Through spinodal decomposition or condensation nuclei, the phase-separated state with non-zero momentum can be achieved. In this scenario, we observe a gap between coexisting phases, arising not only from the variations in energy density, but also from the distinctions in momentum density or longitudinal pressure. These disparities are characterized by flow velocity and latent heat. Furthermore, by introducing an inhomogeneous scalar external source to simulate a fixed obstacle, we reveal the dynamical response of momentum loss in the moving system. Notably, starting from an initial phase-separated state with uniform flow velocity, and subsequently interacting it with an obstacle, we find that the moving high-energy phase exhibits four characteristic dynamical behaviors—rebounding, pinning, passing, and splitting. These behaviors depend on the velocity of the phase and the strength of the obstacle.
I Introduction
@@
The dynamics of strongly coupled quantum fluids is of great research significance in particle physics experiments and astronomical observations, such as the transport properties of the quark gluon-plasma (QGP) [1] and the merger process of binary neutron stars [2]. The challenge is to construct a theoretical framework to describe the real-time evolution of strongly interacting matter. Fortunately, by dualizing a quantum many-body system to a classical gravitational entity with an extra dimension, holography [3, 4, 5, 6] converts such a problem into solving a system of time-dependent gravitational equations, which is always amenable to the numerical relativity [7].
@@
A breakthrough related to holographic fluids is the description of dynamics near a thermal first-order phase transition. Initially, in order to mimic the equation of state of quantum chromodynamics (QCD), the holographic first-order phase transition model was constructed bottom-up [8, 9], involving a real self-interacting scalar field with an external source. Subsequently, the dynamical stability of the thermal phases in equilibrium was revealed through linear perturbation theory [10, 11]. The results indicate the existence of a branch of hydrodynamic modes that dominates the linear stability of the system. For the spinodal region with negative specific heat, the hydrodynamic modes become dynamically unstable in the long wavelength region, similar to Gregory-Laflamme instability [12, 13]. Going a step further, the real-time dynamics of the spinodal instability was numerically simulated at the nonlinear level [14, 15, 16, 17, 18, 19], revealing the second-order hydrodynamic description of the dynamical process and the phase-separated configuration of the final state. On the other hand, for the thermodynamically metastable region adjacent to the spinodal region, the occurrence of a dynamical transition requires the system to cross an unstable critical state [20, 21, 22], in which a critical phenomenon analogous to that in gravitational collapse [23, 24, 25, 26] occurs, indicating a nonlinear instability. Furthermore, through holographic quench, the nonlinear instability of the phase-separated state was also revealed [27], achieving a bidirectional dynamical transition between homogeneous and inhomogeneous configurations. At this point, the theoretical framework on the dynamical stability of the holographic first-order phase transition system has been roughly completed.
@@
Beyond stability analysis, the dynamic properties of holographic fluids near phase transitions and their applications have also been extensively investigated. An interesting research topic is the simulation of cosmological phase transitions [28, 29, 30], with the resulting gravitational wave spectrum potentially observable to detectors [31, 32, 33]. For the case of cosmological phase transitions proceed via bubble nucleation, the dynamic description of the domain wall connecting the nucleated bubble of a stable phase and the overcooled medium was successfully revealed [34, 35, 36, 37], in which the velocity of the domain wall, a crucial parameter on which the gravitational wave spectrum depends, is quantitatively characterized. On the other hand, if the Universe supercools along the metastable branch and enters the spinodal region, there will be a cosmological phase transition induced by the exponential growth of unstable modes, resulting in a qualitatively different gravitational wave spectrum [38]. Another important application lies in the QGP produced in relativistic heavy ion collisions [39], which are viewed as collisions of gravitational shock waves in the dual description [40, 41, 42, 43, 44, 45]. If QCD possesses a critical point, then the far-from-equilibrium dynamics of QGP across a phase transition can be effectively captured by a holographic model with a phase transition [46, 47, 48]. Other related content includes holographic turbulence [49, 50], holographic superfluid [51, 52, 53], etc.
@@
Inspired by the fact that fluids in reality generally have a flow velocity, we will study the real-time dynamics of the holographic steady flow, which is dual to the boosted black brane geometry on the gravity side [54, 55], near a thermal first-order phase transition. The plan of the paper is as follows. In the following section II, we will introduce a holographic first-order phase transition model through the thermodynamic properties of the equilibrium system. In section III, we will study the dynamical transitions of holographic steady flows through fully nonlinear numerical simulations, revealing spontaneous and critical dynamical processes, which correspond to unstable and metastable initial states, respectively. Similarly, the final state is a two-phase coexisting geometric configuration. The difference is that the high-energy phase moves in the low-energy bath at a uniform velocity.
@@
A natural question is whether there is a damping mechanism that slows down the moving domain. We will provide such a mechanism in section IV by introducing an inhomogeneous external source to the scalar field. From the Ward-Takahashi identity, such a local inhomogeneous scalar source can effectively induce changes in the momentum of the system, thereby damping the motion of a steady flow or a moving domain, which is verified by the dynamical results. In particular, we reveal four characteristic dynamical behaviors of the moving domain under interaction with the local potential barrier: rebounding, pinning, passing through, and splitting, depending on the flow velocity of the moving domain and the strength of the potential barrier. Finally, we end this paper with a conclusion and outlook in section V.
II Holographic model
@@
In this section, we will introduce a holographic model with a thermal first-order phase transition from equations of motion, to Ward-Takahashi identity, to thermodynamic phase diagrams involving static states and steady flows.
II.1 Gravity setup
We consider a four-dimensional holographic model involving gravity and a scalar field, described by the following renormalized action [56, 57, 58]
@@ II.1 Gravity setup
According to the AdS/CFT correspondence, such a gravitational system is dual to a conformal field theory deformed with a scalar operator at the three-dimensional boundary. The reason we shy away from working on a five-dimensional holographic model dual to a four-dimensional field theory is to avoid the appearance of holographic anomaly [57], which would induce the field functions to exhibit a logarithmic asymptotic behavior at the infinite boundary, thus greatly reducing the accuracy and convergence of the numerical method. On the other hand, the dynamics of the system depends on the form of the self-interaction of the scalar field [9]. Such self-interaction (3) will cause the phase diagram structure of the model to have a spinodal region, where the states suffer from a long-wavelength instability [10], leading to the occurrence of first-order phase transition dynamics. Such a setup can approximately capture the physical phenomena in the general case. In order to better match QCD, a five-dimensional holographic model should be considered, in which the selection rules of the self-interaction potential are derived from the QCD data. More generally, one can consider the general case of additional electromagnetic fields [59]. These extensions bring certain challenges to numerical techniques.
@@
where all metric fields are functions of $(t,x,r)$, and the translational invariance in the $y$-direction is preserved. In this case, the non-zero components of the energy-momentum tensor (6a) are $\left(\epsilon,\eta,p_{L},p_{T}\right)=\left(-T^{t}_{t},-T^{t}_{x},T^{x}_{x},T^{y}_{y}\right)$, which represent the energy density, momentum density, the longitudinal and transverse pressures of the system, respectively. Since the system we consider is homogeneous in the $y$-direction and has no flow, there is no momentum in the $y$-direction and the shear force, so the corresponding energy-momentum tensor components are all zero $\left(T^{t}_{y},T^{x}_{y}\right)=\left(0,0\right)$. The non-trivial components of the Ward-Takahashi identity (8) reduce to
@@
Based on the second relationship above (10b), we expect that the local inhomogeneous scalar source can provide a damping mechanism for the moving system to evolve from a steady flow to a static state.
II.2 Phase diagram
@@
In order to obtain the steady flow with non-zero momentum density, we need to abandon such an assumption and retain all components of the metric ansatz (9). The system at this time has two independent physical degrees of freedom. Note that in this section we are concerned with the time-independent homogeneous solution, where the field functions depend only on the coordinate $r$, and set the scalar source to $\phi_{0}=1$. After obtaining the geometric configuration and setting the effective Newton constant $\kappa^{2}_{4}=8\pi G_{4}=1$ with the four-dimensional Newton constant $G_{4}$, following black hole thermodynamics, one can extract the temperature and entropy density of the system from the surface gravity and area element of the black hole horizon respectively, as follows
@@
where $r_{h}$ represents the location of the event horizon. For our case without the additional Noether flow, the free energy density satisfies the following thermodynamic relation [60]
@@
Figure 1 shows the phase diagrams under different flow velocities, which are related through the usual boost transformation. For a steady flow with flow velocity $v$, after applying the Lorentz transformation
@@
with a boundary one-form $\omega=\gamma\left(-dt^{\prime}+vdx^{\prime}\right)$. Such a metric form describes a comoving fluid system relative to an observer whose $3$-velocity components equal $\omega^{i}$ [7]. To this moving observer, the fluid is at rest. Note that in the frame (9), the boundary one-form is $-dt$, indicating a stationary observer. Our numerical results show that for any observer, the relatively stationary system has the same phase structure, that is, the case of $v=0$ in figure 1. In other words, in the frame (14), the phase structure of the system does not depend on the parameter $v$.
@@
Recently, simulating cosmological phase transitions via nucleation of bubbles in a holographic first-order phase transition model, the velocity of planar domain walls during nucleation induced by a supercooled state is revealed [34]. The numerical simulation results show that the velocity of the domain wall depends linearly on the pressure difference between the inside and outside of the bubble, namely $\Delta p\sim v$. This is a non-equilibrium situation, in which the bubble is growing. As a complement, for a steady-flowing bubble, the gap (17) indicates that the velocity of the domain wall has a different dependence on the pressure difference, expressed as $\Delta p\sim v^{2}$. Such results indicate that the dependence of domain wall velocity on the pressure difference between the inside and outside of the bubble depends on the state of its motion. On the other hand, the topological structure of the domain wall is also an important factor affecting such dependence [61].
@@
Quantitatively, figure 2 shows the phase transition temperature and latent heat of the first-order phase transition as a function of the flow velocity. The monotonically decreasing curve in figure 2 indicates that a steady flow with a faster flow velocity has a lower phase transition temperature. Such a result shows that for a fast-flowing fluid, the first-order phase transition is difficult to occur. On the other hand, as shown in figure 2, the latent heat of phase transition increases monotonically with the increase of flow velocity, indicating that the faster the fluid flows, the more energy needs to be absorbed or released to transform between different thermal phases. From relationship (17), the increasing latent heat ensures that the pressure difference between the coexisting phases increases with the increase of flow velocity, providing sufficient driving force for the movement of the domain wall. An interesting question is what happens when the flow velocity approaches the speed of light. Unfortunately, due to numerical problems, we cannot obtain the geometric configuration for the case $v>0.8$, leaving this question unresolved and awaiting further exploration in the further. However, from the trend of the curves in Figure 2, we can guess that a possible result is that when the flow velocity approaches the speed of light, the phase transition temperature approaches zero and the latent heat tends to infinity, leading to the disappearance of the first-order phase transition. In order to verify this speculation, a possible strategy is to use an analytical approximation to solve the static field equations to extract the physical quantities in the extreme case.
III Phase separation
@@
III.1 Spinodal decomposition
For the spinodal region where the specific heat is negative, the thermodynamic instability of the system induces a dynamical instability [10, 11]. At the linear level, there is a branch of hydrodynamic modes where the imaginary part of the frequency is positive in the long wavelength region. Under arbitrarily small perturbations, such unstable modes will grow exponentially and push the system away from the initial equilibrium state, inducing a dynamical transition.
In order to reveal such a dynamical process in the case of steady flow, we select a steady flow in the spinodal region as the initial data and then apply a small perturbation. Figure 3 shows the real-time dynamics of the energy density of the system, from which it can be seen that a dynamical transition occurs that destroys the spatial translational symmetry. The entire dynamical process can be roughly divided into three stages. Initially, several energy peaks are spontaneously excited from the homogeneous geometric configuration and move along the flow direction of the background. Subsequently, these energy peaks merge together to form a moving domain, accompanied by fluctuations in energy density. Finally, as the energy fluctuations decay, the system gradually settles down to a phase-separated state with non-zero momentum. During this evolution, from the Ward-Takahashi identity (10), the total energy and total momentum of the system are conserved.
The final state of the evolution consists of coexisting phases at the phase transition temperature, which occupy different spatial regions and are connected by domain walls. The conservation of total momentum results in the system not being at rest, but flowing steadily at a uniform velocity as in the initial state. Since the momentum density of the low-energy phase is close to zero, as shown in figure 1, such a geometric configuration can be approximated as a dynamic structure of a moving domain of the high-energy phase in the background of a static low-energy bath. Note that the velocity of the moving domain in the final state is $v=0.37$, compared with the velocity $v=0.4$ of the initial steady flow, indicating that the phase transition reduces the flow velocity of the system. For such a time-dependent inhomogeneous stationary state with uniform flow velocity $v$, there exists a Killing vector
@@ III.1 Spinodal decomposition
III.2 Condensation nucleus
For the thermodynamically metastable region adjacent to the spinodal region, the system is dynamically stable at the linear level due to the absence of unstable modes, indicating that small perturbations will be dissipated and are not sufficient to change the state of the system. However, inspired by the fact that condensation nuclei induce phase transitions of supercooled water in reality, the thermodynamically metastable state is verified to suffer from a type of nonlinear instability [20, 21, 22]. In this case, the system is in a local ground state. The occurrence of a phase transition requires a disturbance of sufficient intensity to push the system across a dynamical barrier and thereby evolve to another local ground state. Near the threshold of the disturbance intensity, a critical state will appear.